OpenStudy (cutiecomittee123):

Suppose a triangle has two side lengths of 4 and 4 and the angle between the side sis pi/3. What is the length of the third side of the triangle?

OpenStudy (cutiecomittee123):

I used the SAS formula and came out with 3 as the answer

OpenStudy (caozeyuan):

it is 4 because pi is 180 and so pi/3 is 60, 180-60=120, since 4=4, two angles are 120/2=60, so all three sides are same

OpenStudy (victoriasushchik):

Use Law of Cosines: Let a,b,c be the side lengths and let C be the angle between a and b. Then c^2 = a^2 + b^2 - 2ab cos C Now plug in a = 2, b = 3, C = pi/3 c^2 = 2^2 + 3^2 - 2(2)(3) cos (pi/3) c^2 = 4 + 9 - 12*(1/2) c^2 = 4 + 9 - 6 c^2 = 7

OpenStudy (caozeyuan):

@Victoriasushchik WHERE are the 2 and 3 coming from

OpenStudy (victoriasushchik):

thats your question i just answered it t